OpenStudy (anonymous):

solve 3/x-2+1=-23+7x/(x-2)(x-3)

OpenStudy (anonymous):

are there any parenthesis in the first part or is it just x-2+1?

OpenStudy (anonymous):

yea that IS a bit confusing.... :P

OpenStudy (anonymous):

the way you write the question makes it a bit hard to understand

OpenStudy (anonymous):

wat u talkin about

OpenStudy (anonymous):

u need to space it out a bit that would make it easier to solve for both us and urself :)

OpenStudy (anonymous):

3 / x-2 +1 = -23+7 / (x-2) (x-3)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im waiting on u to solve it

OpenStudy (anonymous):

Does it look like \[3/(x-2)+1=-23+7x/(x-2)(x-3)\] ?

OpenStudy (anonymous):

\[(3+x-2)(x-3)=-23(x-3)+7x\] \[(1+x)(x-3)=-23x+69+7x\]

OpenStudy (anonymous):

Now you can do the rest, can't you coreeta?

OpenStudy (anonymous):

can u solve x for me

OpenStudy (anonymous):

Well, that would be as follows \[x-3+x^2-3x=-16x+69\]

OpenStudy (anonymous):

\[x^2+14x-72=0\] If I am not mistaken anywhere through out the solution

OpenStudy (anonymous):

Do you know how to solve quadratics equation?

OpenStudy (anonymous):

\[{{x = (31 -Sqrt[151])/12}, {x = (31 + Sqrt[151])/12}} \]

OpenStudy (anonymous):

what is the values of x in this problem

OpenStudy (anonymous):

I think I have done it wrong somewhere. I mean some silly mistake It must be \[3/x+7/(x-2)+15 = 0\]

OpenStudy (anonymous):

\[3/x = (23-15 x)/(x-2)\]

OpenStudy (anonymous):

Now you got to cross multiply, and then solve the quadratic equation

OpenStudy (anonymous):

i dont no know to do that im homeskooled

OpenStudy (anonymous):

Which level you are currently in?

OpenStudy (anonymous):

im n 9th . can u plz find the valuse of x for me

OpenStudy (anonymous):

Do you just need the answers or you need to understand the steps?

OpenStudy (anonymous):

i need the answer

OpenStudy (anonymous):

Okay, then forget everything I did there. I did several mistakes. Actually its quite tough to type things here

OpenStudy (anonymous):

I am doing it on a piece of paper, and telling you the answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

pi when you finish tell me if ur answer look anything like this x=-0.252 and x=1.585

OpenStudy (anonymous):

First coreeta, have a look at this, Does your equation looks like the one I have attached

OpenStudy (anonymous):

If so then your solutions are

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

and the another solution for x is

OpenStudy (anonymous):

These are the final answers and they are correct, (I am 100% sure)

OpenStudy (anonymous):

The solutions are \[x = 1/12 (31-\sqrt(151))\] \[x = 1/12 (31+\sqrt(151))\]

OpenStudy (anonymous):

Is it okay coreeta?

OpenStudy (anonymous):

yes